OpenStudy (abmon98):

1-tan^2(theta)/1+tan^2(theta)=1-2sin^2(theta) Prove identity

Parth (parthkohli):

Do you know a common Pythagorean identity associated with this?

OpenStudy (abmon98):

Sin^2(theta)+Cos^2(theta)=1

OpenStudy (abmon98):

tan^2(theta)=Sin^2(theta)/cos^2(theta)

Parth (parthkohli):

well, you can actually just use the second thing. forget identities.

OpenStudy (abmon98):

i used it and still cant reach to 1-2sin^2(theta)

Parth (parthkohli):

what are you getting?

OpenStudy (abmon98):

1-(Sin^2(theta)/cos^2(theta))/1+(sin^2(theta)/Cos^2(theta)

Parth (parthkohli):

Get a common denominator in both the numerator and denominator.

Parth (parthkohli):

Then cancel that denominator.

OpenStudy (abmon98):

How to get a common denominator and they are already similar but the sign are different

OpenStudy (***[isuru]***):

@ParthKohli it would be easy if u simplify both R.H.S and L.H.S and go to cos 2x .. R.H.S is already in the expanded form of cos 2x and under 3 steps u could reach cos^x - sin^x from L.H.S which is also cos 2x :)

Parth (parthkohli):

\[1 - \tan^2\theta = 1 - \dfrac{\sin^2\theta}{\cos^2 \theta} = \frac{\cos^2\theta - \sin^2\theta}{\cos^2\theta}\]

Parth (parthkohli):

similarly in the denominator, you have\[\frac{\sin^2\theta + \cos^2\theta}{\cos^2\theta}\]Cancel the \(\cos^2\theta\)

Parth (parthkohli):

also, \(\sin^2\theta + \cos^2 \theta = 1\).

Parth (parthkohli):

So\[\dfrac{\cos^2\theta- \sin^2\theta}{\sin^2 \theta + \cos^2\theta} = \cos^2\theta - \sin^2\theta\]

Parth (parthkohli):

now in \(\cos^2\theta - \sin^2\theta\), substitute \(1 - \sin^2\theta\) for \(\cos^2\theta\)

OpenStudy (abmon98):

Thank you for your help :)

Parth (parthkohli):

You're welcome!